OpenStudy (anonymous):

Why is it reasonable for the value of “any number’ to decrease when it is multiplied by 0.01? Justify your response.

OpenStudy (anonymous):

Because of how small 0.01 is

OpenStudy (anonymous):

2 times 0.01 is 0.02

OpenStudy (anonymous):

1000 times 0.01 is 10 which is considerably smaller

OpenStudy (anonymous):

When you have such a number so close to zero these things happen

OpenStudy (anonymous):

yea but when you think of multiplication you think the number is going to increase, how come it decreases when multiplied by 0.01 74.2 x 0.01=0.742 74.2/0.01=7420

geerky42 (geerky42):

According to Multiplicative Identity, multiplying number by \(1\) will leave it unchanged. \(n\times1=n\) So multiplying a number \(n\) by any number between \(0\) to \(1\) will make result smaller than \(n\). Likewise, multiplying a number \(n\) by any number greater than \(1\) will make result larger than \(n\) Does that help?

OpenStudy (anonymous):

yes, thank you its more clear if you had to explain it with place value and multiplying and dividing by powers of ten how can this be explained?

geerky42 (geerky42):

Maybe say something like that decimal can be convert into fraction, you can see that for example: \(0.231 = \dfrac{231}{1000}\); it will make n smaller since \(1000>231\) Dividing number \(n\) by any number \(m\) is same as multiplying n by reciprocal of m \(n\div m = n\times\dfrac{1}{m}=\dfrac{n}{m}\) I am not sure exactly how to explain this, but what I said should help you, right?

OpenStudy (anonymous):

right, ty so much again

OpenStudy (anonymous):

isn't 0.01 in fraction 100?

OpenStudy (anonymous):

in fraction is 1/100

OpenStudy (anonymous):

kk,why does greeky42 231/1000